# On the Utility of Domain-Adjacent Fine-Tuned Model Ensembles for Few-shot Problems

### v1

@@ Abstract
Large Language Models (LLMs) have been observed to perform well on a wide range of downstream tasks when fine-tuned on domain-specific data. However, such data may not be readily available in many applications, motivating zero-shot or few-shot approaches using domain-adjacent models. While several fine-tuned models for various tasks are available, finding an appropriate domain-adjacent model for a given task is often not straight forward. In this paper, we study DAFT-E, a framework that utilizes an Ensemble of Domain-Adjacent Fine-Tuned Foundation Models for few-shot problems. We show that for zero-shot problems, this ensembling method provides an accuracy performance close to that of the single best model. With few-shot problems, this performance improves further, at which point DAFT-E can outperform any single domain-adjacent model while requiring much less data for domain-specific fine-tuning.
On the Utility of Domain-Adjacent Fine-Tuned Model Ensembles for Few-shot Problems
Md Ibrahim Ibne Alam and Koushik Kar Rensselaer Polytechnic Institute NY, USA, 12180                        Parikshit Ram, Soham Dan and Horst Samulowitz IBM, Yorktown Heights NY, USA, 10598
@@ 1 Introduction
Pre-trained Large Language Models (LLMs), are used for different downstream tasks. Usually, for LLMs that are not directly suitable for downstream tasks, a task-specific header layer is needed at the output. We define the base model as the LLM with a task-specific header layer. As an example, BERT models Kenton and Toutanova (2019) were trained for sentence completion tasks. In order to perform sentiment classification, we need to add a header layer with output dimension matching the number of classes. Usually, base models do not need much training to perform well, because the pre-trained LLMs have already been trained on a large corpus of data. To perform a task with base models, we can fine-tune a base model with task-specific training data, and then use the fine-tuned model. However, it is often the case that appropriate training data for fine-tuning is not readily or abundantly available. Moreover, even if one does have the data to fine-tune, one may not have the computational resources, or the time required, to fine-tune these base models. It is worth noting that fine-tuning can be computation, memory and time intensive, even if very few iterations are needed to attain good performance. Thus, we explore alternatives that bypass those challenges and leverage the large number of fine-tuned LLMs that are already available in curated repositories such as Hugging-face HF (2024c); Kim (2023). It is important to acknowledge the scale of fine-tuned models being created. For instance, just three days after LLaMA-3 was released, more than 1,000 fine-tuned LLMs were publicly available on Hugging-face HF (2024a).
@@
Here we show the potential of these existing and publicly available fine-tuned models (e.g., including those fine-tuned via LORA Sheng et al. (2023) and PEFT Li and Liang (2021)) by leveraging them to perform different tasks. To use a fine-tuned model on a task, e.g., sentiment classification, we need that model to be domain adjacent (formally defined and discussed in Section 3). Informally, a fine-tuned LLM is domain-adjacent if (i) it has been fine-tuned on a similar general task, such as sentiment analysis, and (ii) its output space can be appropriately mapped to the current (few-shot) task. We denote these Domain Adjacent Fine-Tuned models as DAFT. In the generation process of DAFT, the closer the fine-tuning dataset is to the target domain data, the better will be the performance of DAFT. One main property of DAFT is that it can be used without any further training (no extra computation) in a zero-shot manner (no data adaptation). To perform inference with DAFT we do not require large computational resources compared to when performing fine-tuning. It is important to note that when using DAFT, a method to choose among available model(s) to achieve improved performance for a given task is required (see Section 4).
The availability of DAFT models combined with their ease of use, make them very appealing to use for different tasks. To analyze the efficiency of using these DAFT models in comparison to other existing options, we first define several comparable alternatives. Let us denote $D_{T}$ as the testing dataset on which we need to perform a given task. Then we have the following options (Figure 1):
@@
    DAFT ${}^{\text{Z}}$ – Pick one available DAFT model and use it for zero-shot inference (no training cost).
@@
    DAFT-E ${}^{\text{Z}}$ – Use a subset or all DAFT models and ensemble them for zero-shot inference.
@@
    DAFT-E – Use a subset or all DAFT models and perform few-shot learning of ensemble weights using training data from $D_{T}$ for inference.
Figure 1: Different options for few-shot tasks given LLMs and community created DAFT models: (1) Directly fine-tune the LLM as FT; (2) select one DAFT model and use it zero-shot as DAFT ${}^{\text{Z}}$, or (3) with task-specific fine-tuning as DA(FT)2. (4) DAFT models can be ensembled and used zero-shot as DAFT-E ${}^{\text{Z}}$, or (5) with extremely lightweight few-shot adaptation using weighted ensembling as DAFT-E.
In the next section we discuss related work, followed by Section 3 defining DAFT models and evaluating their performance on various NLP tasks. Section 4 discusses zero-shot and few-shot performance, while Section 5 provides some theoretical performance bounds for the ensembling methods.
@@ 2 Related Work
Ensembling is a technique that combines the predictions of multiple classifiers to generate a single decision and is extensively investigated Breiman (1996); Clemen (1989); Maclin and Opitz (1997); Dietterich (2000). The pre-requisites for effective ensembling are: (i) models with decent performance, and (ii) models that make independent errors, i.e., training diversity in models Ovadia et al. (2019); Lakshminarayanan et al. (2017). The ensemble of models can be the weighted or unweighted average of the model outputs. For weighted averaging, some training is needed to learn the weights Caruana et al. (2004). Isotropic merging Polyak and Juditsky (1992) is another technique similar to ensembling that has been studied alongside ensemble methods over the last few decades and has shown promising performance Wortsman et al. (2022). The main novelty of our work is the use of ensembling in the context of DAFT models fine-tuned from pretrained LLMs, to maximize the efficacy of using these models.
Another approach related to ensembling is blending of models. For instance, LLM-Blender Jiang et al. (2023) performs ensembling of $n$ LLMs by pairwise ranking and generative fusion. The pairwise ranking approach requires creating a custom dataset, training a BERT model, and it incurs substantial computational overhead when performing inference. The proposed generative fusion combines the ranked list from the pairwise ranking approach with a fine-tuned LLM for this task to generate a response. In contrast, our approaches focus on a completely task agnostic (e.g., not just generative tasks) and computationally low-cost solution to leverage multiple fine-tuned LLMs. Another type of blending is introduced in Lu et al. (2024), and it is orthogonal to our approach since it selects base models at random and combines them by adding the response of an already evaluated model to the input of the subsequent model. While this knowledge sharing makes it a sequential approach by design, it can be readily combined with ours.
Mixture of Experts (MoE) Jacobs et al. (1991) has been utilized by different ML models for both regression and classification tasks Yuksel et al. (2012). Before recent developments in LLMs, MoE was mainly focused on dense expert models. However, that idea has been transformed into sparse expert models with the introduction of MoE in LLMs. In current practice, MoE is a sparse expert model, where a set of parameters of a neural network are partitioned into experts having their own weights Fedus et al. (2022). One recent work on MoE, Mixtral Jiang et al. (2024) uses a sparse LLM MoE where a routing network (e.g., Rosenbaum et al. (2017)) is pre-trained to map input tokens to a subset of experts. Compared to all these MoE models, our approach does not require intensive pre-training, and the most computationally expensive variant that uses weight-learning (DAFT-E) is only trained in a few-shot fashion based on input examples.
@@
Transfer learning Pan and Yang (2009); Zhuang et al. (2020) is closely related to the goals of our work here, where we wish to adapt a model trained on a data-rich source distribution to a data-scarce target problem. The success on the target problem often relies heavily on the level of positive transfer one can achieve, and various ways of training the model on the source data and the target data have been developed to maximize the positive transfer. In the context of LLMs, fine-tuning a (base) LLM with task-specific data is an instance of transfer learning. In our paper, we consider the DAFT models as the source models. However, not every DAFT model may transfer positively, since the DAFT models are obtained from public non-curated pools.
@@ 3 DAFT Models
With the rapid expansion of open platforms like Colab and Kaggle to perform model training, and having access to different public datasets, researchers can access LLMs and datasets that can be used to fine-tune these LLMs. This has lead to a large repository of publicly available fine-tuned base models (as defined in Section 1) HF (2024c); Kim (2023). In Section 1 we argued that the performance of any such fine-tuned model on some target dataset depends on the source dataset on which it was fine-tuned on. We define a dataset to be domain adjacent dataset if it can be used for the same (general) task as the target dataset. We call a model a domain adjacent fine-tuned model (DAFT) if: (i) it was fine-tuned with a domain adjacent dataset, and (ii) the output space of the model can be appropriately mapped to the target task.
@@ 3.1 Formation of DAFT models
3.1.1 Datasets
We consider two broad NLP tasks with datasets from different sources HF (2024b); University (2024); Kaggle (2024), for (1) Sentiment Analysis (positive or negative sentiment): Amazon polarity, Cornell Movie, IMDB, SST2, Tweet sentiment, and Yelp Polarity; and (2) Textual Similarity (if two sentences are similar): MRPC, QQP, STS-B (details in Appendix A.1).
@@ 3.1.3 Base Models
A base model is created by adding a header layer to an LLM. Since the base models are not trained to perform any specific task and the header layer has not been fine-tuned, the performance of all the base models is similar to random guessing. For a specific dataset we can fine-tune the base models, and depending on the extent of fine-tuning, we obtain a partial (few-shot) or fully fine-tuned model. For some base model $B$ and dataset $D$, let us denote $\Phi(B,D,n)$ as the fine-tuned version of $B$ on $n$ amount of data from dataset $D$. If the parameter $n$ is missing as the argument, the base model is assumed to be fully fine-tuned (FFT) on $D$. In our experiments, with few shot fine-tuning, we vary $n$ in the range of $2-256$ samples. For the (FFT) models, we fine-tune until loss stabilization [^2] We fine-tune until the loss fluctuation of less than ±$\pm 1\%$..
@@ 3.2 Performance of DAFT Models
Figure 2: Efficacy of DAFT models: The vertical axis corresponds to the data used to fine-tune a base LLM (here we show Roberta-base; the heatmaps for the other models show a similar trend), and the horizontal axis corresponds to the unseen test data. All datasets have matching output spaces and pertain to a sentiment classification problem. The diagonal is the performance of FT– the LLM tested with data from the distribution it was fine-tuned with , denoting the upper watermark for that test data (only exception being SST2 – discussed in text). The off-diagonal entries correspond to the performance of the DAFT– LLM trained on one dataset, and tested on another, highlighting the potential computationally cheap (free!) zero-shot benefit that DAFT models can provide. Note that, for any given test data, the performance of a DAFT LLM can vary significantly.
@@
Let us denote the train and test splits of the target dataset as $D_{T}^{\prime}$ and $D_{T}^{\prime\prime}$ respectively. Since the standard train and test splits are identically distributed (iid), we expect a model trained on the target data ($D_{T}^{\prime}$) to perform the best on the test data $D_{T}^{\prime\prime}$. This performance value is the ceiling benchmark that we aim to match or surpass. For that purpose, we performed full fine-tuning (FFT) on all three base models using training data of all datasets of sentiment analysis and textual similarity, to come up with 18 FFT models for sentiment analysis and 9 FFT models for textual similarity. Hence, for any target dataset, we can have 15 DAFT models for sentiment analysis by excluding the three models that were fully fine-tuned on $D_{T}$, and 6 DAFT models for textual similarity task.
@@
We plot the individual heat maps with the FFT models on sentiment analysis and textual similarity tasks in Figure 2. It is evident from the figure that the performance of the Cornell and Tweet datasets do not match the other four datasets’ performance on the sentiment analysis task. Our empirical study suggests that the test data of Cornell is distributionally different from the other datasets, and for all FFT models the performance was poor. The Tweet dataset seems to have a different distribution in both its training and testing data (compared to the other 5 datasets). Looking at the Tweet row and column, we observe that apart from the diagonal element, all the other accuracies are small, and a high $94\%$ accuracy is attained when the model is trained and tested on Tweet data. Upon inspecting the row of SST2, another interesting observation can be made, the performance of DAFT-Cornell beats that of DAFT-SST2. This is caused by the test dataset of SST2 being more aligned with the training data of Cornell, with further details in Appendix A.3.1. Lastly, for textual similarity datasets, the performance of DAFT models was not as good as the sentiment analysis, but still show promise.
@@
[linecolor=black!0,backgroundcolor=orange!10]These results show that DAFT models can potentially provide very competitive zero-shot performance on new tasks.
3.2.1 Performance of DAFT ${}^{\text{Z}}$
Figure 3: Relative Improvement of DAFTs compared to the Single Best DAFT for sentiment analysis: For each test dataset, we consider 15 DAFT LLMs (fine-tuned on data different from the test data), and show the RI of DAFT ${}^{\text{Z}}$ over the single-best DAFT LLM (out of the 15) (the colored circles $\bullet$). Values less than 0 indicate performance degradation. The red squares $\blacksquare$ denote the performance of zero-shot prompting on larger LLMs (Bart-large-mnli, Roberta-large-mnli, and Opt-1.3b).
Figure 3 shows the relative improvement $(RI)$ of all the DAFT models for sentiment analysis task when compared to the single best performing DAFT model. [^3] Thus, all the $RI$ values (corresponding to the other DAFTs) will be non-positive (zero for the best single DAFT). The relative improvement $(RI)$ of all the DAFT models for textual similarity task is provided in Appendix (Figure 14). From Figure 3 it is evident that there are some DAFT models that are not as good as the single best and for some specific choices (as for Tweet), the performance can be poor. On the other hand, among the three larger models (models from $\mathcal{L}$), one of them (opt-1.3b) showed poor performance on most of the datasets, while the other two performed quite well. However, these larger models still fails to beat the single best DAFT in all cases, except Tweet. This low performance of DAFT models on Tweet is potentially because of the difference in the nature of this dataset. {mdframed}[linecolor=black!0,backgroundcolor=orange!10]These results show that zero-shot performance of DAFT models can vary significantly, making the choice of the DAFT model critical.
3.2.2 Comparison of FT and DA(FT)2
Table 1: Average Relative Improvement of DA(FT)2 compared to FT (Top: 6 sentiment analysis tasks. Bottom: 3 textual similarity tasks): Positive values show DA(FT)2 improvement over FT; larger values imply higher improvements. Please see Table 6 and 7 in Appendix for detailed results. We see that DA(FT)2 significantly improves over FT in few-shot problems.
Dataset 2-shot 4-shot 8-shot 16-shot 32-shot 64-shot 128-shot

### v2

# On the Utility of Existing Fine-Tuned Models on Data-Scarce Domains

@@ Abstract
Large Language Models (LLMs) have been observed to perform well on a wide range of downstream tasks when fine-tuned on domain-specific data. However, such data may not be readily available in many applications, motivating zero-shot or few-shot approaches using existing domain or task adjacent (fine-tuned) models, which we call DAFT. While several fine-tuned models for various tasks are available, finding one appropriate DAFT model for a given task is often not straight forward. In this paper, we explore different utilization techniques of these existing DAFT models for data-scarce problems, i.e., tasks for which data is not available or limited. We observe that for zero-shot problems, ensembling of DAFT models provides an accuracy performance close to that of the single best model. With few-shot problems (few data from target domain available), this performance can be improved further by picking or putting more weights to the DAFT models that are expected to perform better on the target task.
@@ 1 Introduction
Pre-trained Large Language Models (LLMs) are used for different downstream tasks. Usually, for LLMs that are not directly suitable for downstream tasks, a task-specific header layer, e.g., for classification, is needed at the output. We define the base model as the pre-trained LLM with a task-specific (un-tuned) header layer. As an example, BERT models (Kenton & Toutanova, 2019) were trained for sentence completion tasks. In order to perform sentiment classification, we need to add a header layer with output dimension matching the number of classes. Usually, base models do not need much training to perform well, because the pre-trained LLMs are already trained on a large corpus of data.
To perform a task with base models, we can fine-tune a base model with task-specific training data, and then use the fine-tuned model. This fine-tuning of model is computationally much less extensive compared to the initial training phase of the LLMs. However, it is often the case that appropriate training data for fine-tuning is not readily or abundantly available, i.e., the domain is data-scarce. Moreover, even if one does have the data to fine-tune, one may not have the computational resources, or the time required, to fine-tune. It is worth noting that fine-tuning can be computation, memory and time intensive, even if very few iterations are needed to attain good performance. Hence, we compare different fine-tuning solutions with some alternatives that bypass the challenges of fine-tuning and leverage the large number of fine-tuned LLMs already available (in curated repositories like Hugging-face (HF, 2024c; Kim, 2023)). It is important to acknowledge the scale of fine-tuned models being created; for instance, just three days after LLaMA-3 was released, more than 1,000 fine-tuned LLMs of LLaMA-3 were publicly available on Hugging-face (HF, 2024b).
{mdframed}
[linecolor=black!0,backgroundcolor=orange!10]Research Question: Can we use these plethora of fine-tuned LLMs for a data-scarce task?
Figure 1: Different options for few-shot tasks given LLMs and community created DAFT models: (1) Directly fine-tune the LLM (FT); (2) Select one DAFT model and use it zero-shot (DAFT ${}^{\text{\!\scriptsize Z}}$); (3) Use a DAFT model but with task-specific fine-tuning (DA(FT)2); (4) Ensemble multiple  DAFT models and use it zero-shot (DAFT-E ${}^{\text{\!\scriptsize Z}}$); (5) Use ensemble of DAFT models with extremely lightweight few-shot adaptation of the ensemble weights (DAFT-E).
@@
Here, we investigate the potential of existing and publicly available fine-tuned models (e.g., including those fine-tuned via LoRA (Sheng et al., 2023) and PEFT (Li & Liang, 2021)) by leveraging them to perform different tasks. To use a fine-tuned model on a task, e.g., sentiment classification, we need that model to be domain or task adjacent (discussed in Section 3). We denote these Domain Adjacent Fine-Tuned models as DAFT. One main property of DAFT is that it can be used without any further fine-tuning (no extra computation) in a zero-shot manner (no data adaptation). Also, inference with DAFT does not require large computational resources when compared to performing fine-tuning.
The availability of DAFT models combined with their ease of use, make them very appealing to use for different tasks. However, the performance of DAFT models on the target domain data (denoted as $D_{T}$) can be poor depending on the fine-tuning dataset of the respective DAFT models. In this work, we systematically study four different ways (Fig. 1) of leveraging the DAFT models, and demonstrate how it is possible to get strong performance with low compute adaptation, demonstrating the utility of these community generated DAFT models. These four options along with fine-tuning the base model using data from target domain (FT in Fig. 1) are defined as follows:
@@
    DAFT ${}^{\text{\!\scriptsize Z}}$ – Pick one available DAFT model and use it for zero-shot inference (no training cost).
@@
    DAFT-E ${}^{\text{\!\scriptsize Z}}$ – Use a subset or all DAFT models and ensemble them for zero-shot inference.
@@
    DAFT-E – Use a subset or all DAFT models for few-shot learning of ensemble weights using training data from $D_{T}$ and perform inference.
Section 2 discusses related work, followed by Section 3 discussing DAFT models and their potential for various NLP tasks. Section 4 discusses zero-shot and few-shot performance, while Section 5 theoretically characterizes conditions that result in strong performance from the DAFT ensemble.
@@ 2 Related Work
The two main aspects of our work are: (i) trying to adapt to data-scarce tasks or transfer learning, (ii) leveraging multiple pre-fine-tuned models via ensembling. Here we will briefly discuss related work on transfer learning, ensembling and the closely related blending and mixture-of-experts.
@@
Transfer learning (Pan & Yang, 2009; Zhuang et al., 2020) is closely related to the goals of our work, as we wish to adapt a model trained on a data-rich source distribution to a data-scarce target problem. The success on the target task often relies heavily on the level of positive transfer one can achieve, and various ways of training the model on the source and target data have been developed to maximize the positive transfer (e.g., by extending the already expensive pre-training phase (Gururangan et al., 2020)). A thorough empirical study on the factors for good transfer learning across diverse domains and tasks has been done in Mensink et al. (2021). In Zamir et al. (2018), a parameterized readout function represented by a shallow network was introduced to check (after necessary training) the affinity between any two tasks. Given some budget constraints, using the affinity matrix, the performance on $T$ tasks can be optimized by only training models on $S$ tasks, where $T\gg S$. Here, we consider the DAFT models as the source models for knowledge transfer. However, not every DAFT model may transfer positively, since the DAFT models are obtained from public non-curated pools. Also, for few shot problems, fine-tuning DAFT with target domain data (denoted as DA(FT)2) is a transfer learning option that is considered here (DA(FT)2 in Fig. 1).
Ensembling is a technique that combines the predictions of multiple classifiers to generate a single decision and has been extensively investigated over the past few decades (Breiman, 1996; Clemen, 1989; Maclin & Opitz, 1997; Dietterich, 2000). The pre-requisites for effective ensembling are: (i) models with decent performance, and (ii) models that make independent errors, i.e., training diversity in models (Ovadia et al., 2019; Lakshminarayanan et al., 2017). The ensemble of models can be the weighted or unweighted average of the model outputs. For weighted averaging, some training is needed to learn the weights (Caruana et al., 2004). As another approach, Dvornik et al. (2020) proposed a method to select a weighted concatenation of the output features of models corresponding to multiple domains (models fine-tuned with different datasets) in a few shot environment. The optimized weights of the features are calculated by minimizing the negative log-likelihood classification loss on the few-shot target data. Isotropic merging (Polyak & Juditsky, 1992) is a technique similar to ensembling that has been studied alongside ensemble methods over the last few decades and has shown promising performance with LLMs, e.g., Model Soup and its variants (Wortsman et al., 2022). Recently, a variation of Isotropic merging has been introduced as AdaMerging (Yang et al., 2023), where multiple LLMs, individually expert on different tasks, are merged to generate a single model that can perform well on all tasks. A forward pass on some unlabeled data is needed to find the merging coefficients using entropy minimization. One of the main contributions of our work is to focus on understanding the efficacy of ensembling in the context of DAFT models for data-scarce tasks.
The main differences between Model Soup (Wortsman et al., 2022), AdaMerging (Yang et al., 2023), and the proposed ensemble methods (DAFT-E ${}^{\text{\!\scriptsize Z}}$ and DAFT-E), are highlighted in Table 1. In the table the following abbreviations are used; SLI: Single LLM inference, NMWA: No (LLM model) weight access, MAE: multi-architecture ensemble, SFPA: single forward pass (for few-shot) adaptation. The table highlights that DAFT-E ${}^{\text{\!\scriptsize Z}}$ and DAFT-E provide lightweight LLM backprop-free model combination schemes (similar to Uniform Soup, Greedy Soup and AdaMerging), without requiring access to the model weights, and having the ability to combine LLMs with different architectures. Furthermore, the few-shot adaptation in DAFT-E requires only a single forward-pass with the few-shot samples compared to Greedy Soup, which requires multiple forward-passes. However, both Uniform and Greedy Soup finally need to perform inference with a single LLM while DAFT-E ${}^{\text{\!\scriptsize Z}}$ and DAFT-E have to perform inference with multiple LLMs.
Another approach related to ensembling is blending of models. For instance, LLM-Blender (Jiang et al., 2023) performs ensemble of $n$ LLMs by pairwise ranking and generative fusion. The pairwise ranking approach requires creating a custom dataset, training a BERT model, and it incurs substantial computational overhead to perform inference. The proposed generative fusion combines the ranked list from the pairwise ranking with a fine-tuned LLM to generate a response. In contrast, our approaches with DAFT focus on a completely task agnostic (e.g., not just generative tasks) and computationally low-cost solution to leverage multiple fine-tuned LLMs (Please check Table 1 for comparison). Another type of blending is introduced in Lu et al. (2024), and is orthogonal to ensemble. The method selects base models at random and combines them by adding the response of an already evaluated model to the input of the subsequent model; a sequential approach that can be readily combined with ensembling.
Table 1: Comparison of different methods to combine pretrained models in terms of the features. BPF: Method is LLM backpropagation free. SLI: Method needs to perform inference on the target task with a single LLM. NMWA: No LLM model weight access is required by the method. MAE: Method can handle multi-architecture ensemble. SFPA: Method requires a single forward pass for few-shot adaptation.
Method 0-shot BPF SLI NMWA MAE SFPA
DAFT N/A N/A
Uniform Soup N/A
AdaMerging N/A
DAFT-E ${}^{\text{\!\scriptsize Z}}$ N/A
FT N/A
DA(FT)2 N/A
Greedy Soup
LLM-Blender
DAFT-E
Mixture of Experts (MoE) (Jacobs et al., 1991) has been utilized by different ML models for both regression and classification tasks (Yuksel et al., 2012). Before recent developments in LLMs, MoE was mainly focused on dense mixture of experts, which has been replaced by a sparse mixture of experts (Fedus et al., 2022). One recent work on MoE, Mixtral (Jiang et al., 2024) uses a sparse LLM MoE where a routing network (e.g., Rosenbaum et al. (2017)) is pre-trained to map input tokens to a subset of experts. Compared to all these MoE models (which requires multiple back-propagation through all the involved LLMs to train the routing network), the ensemble of DAFT models that we consider require at most a single forward pass through the LLMs for few-shot target task adaptation, and hence are computationally significantly more lightweight.
@@ 3 DAFT Models
With the rapid expansion of open platforms like Colab and Kaggle to perform model training, and having access to different public datasets, researchers can access LLMs and datasets that can be used to fine-tune these LLMs. This has lead to a large repository of publicly available fine-tuned base models (as defined in Section 1) (HF, 2024c; Kim, 2023). We formally define base model as;
Definition 1.
A base model has a pre-trained LLM and a task-specific (un-tuned) header layer added on top of the LLM.
In Section 1 we argued that the performance of any such fine-tuned base model on some target dataset depends on the source dataset on which it was fine-tuned on. A base model’s header layer can vary depending on the task, and it can become a domain adjacent fine-tuned model (DAFT) for a task when the base model is fine-tuned (either the header layer or the full model) with a dataset that has the same task criterion as the target data. So, we define the DAFT model as follows:
Definition 2.
A domain adjacent fine-tuned model (DAFT) is a fine-tuned base model, if the model can be used for the same (general) task as the target task.
The task similarity between two tasks can be calculated in two ways: (i) by the source and target dataset similarity and (ii) by task description similarity. Since in most cases the source dataset, with which the DAFT model was fine-tuned, is not available, method (ii) is more general. For method (ii), the task similarity can be identified manually or measured using an LLM fine-tuned for ‘textual similarity’. In our definition of DAFT, we are proposing to follow method (ii) to find DAFT models, i.e., when a model is fine-tuned with a dataset that has the same task criterion as the target data.
@@ 3.1 Formation of DAFT models
3.1.1 Tasks and Datasets
We consider two broad NLP tasks, (1) Sentiment Analysis (positive or negative sentiment) and 2) Textual Similarity (if two sentences are similar). Although there are (currently) $1667$ and $936$ DAFT models available for sentiment analysis and textual similarity task in Hugging Face platform, to mitigate the risk of benchmark data leakage, we chose to create our own DAFT models by fine-tuning the base models with specific datasets [^1] The performance of DAFT models downloaded directly from Huggingface are analyzed further in Appendix A.8. (HF, 2024a; University, 2024; Kaggle, 2024). The datasets for Sentiment Analysis are: Amazon polarity, Cornell Movie, IMDB, SST2, Tweet sentiment, and Yelp Polarity; and for Textual Similarity: MRPC, QQP, STS-B (details in Appendix A.1).
@@ 3.1.2 Models
Figure 2: a) LEEP score heatmap: LEEP scores signifying the transferability of knowledge from one dataset to another for sentiment analysis datasets. b) Efficacy of DAFT models: The vertical axis corresponds to the data used to fine-tune a base LLM (here we show Roberta-base; the heatmaps for the other models show a similar trend and are provided in appendix), and the horizontal axis corresponds to the unseen test data. All datasets, for each of the two tasks, have matching output spaces respectively and pertain to sentiment classification and textual similarity problem respectively (left and right). The diagonal is the performance of FT– the LLM tested with data from the distribution it was fine-tuned with, denoting the upper watermark for that test data (only exception being SST2 – discussed in text). The off-diagonal entries correspond to the performance of the DAFT– LLM trained on one dataset, and tested on another, highlighting the potential computationally cheap (free!) zero-shot benefit that DAFT models can provide. Note that, for any given test data, the performance of a DAFT LLM can vary significantly. Also, LEEP score heatmap and DAFT models performance shows very similar trend.
@@ 3.1.3 Base Models
A base model is created by adding a header layer to an LLM. Since the base models are not trained to perform any specific task and the header layer has not been fine-tuned, the performance of all the base models is similar to random guessing. For some base model $B$ and dataset $D$, let us denote $\Phi(B,D,n)$ as the fine-tuned version of $B$ on $n$ amount of data from dataset $D$. If the parameter $n$ is missing as the argument, the base model is assumed to be fully fine-tuned (FFT) on $D$. In our experiments, with few shot fine-tuning, we vary $n$ in the range of $2-256$ samples. For the (FFT) models, we fine-tune until loss stabilization [^3] Fine-tune until the fluctuation of loss is less than ±$\pm 1\%$..
@@ 3.2 Performance of DAFT Models
Figure 3: $RI$ of DAFTs compared to the Single Best DAFT for sentiment analysis: For each test dataset, we consider 15 DAFT LLMs (fine-tuned on data different from the test data), and show the RI of DAFT ${}^{\text{\!\scriptsize Z}}$ over the single-best DAFT LLM (out of the 15) (the colored circles $\bullet$). Values less than 0 indicate performance degradation. The red squares $\blacksquare$ denote the performance of zero-shot classification with larger LLMs, i.e., Bart-large-mnli, Roberta-large-mnli, and prompting with Opt-1.3b.
@@
Let us denote the train and test splits of the target dataset as $D_{T}^{\prime}$ and $D_{T}^{\prime\prime}$ respectively. Since the standard train and test splits are identically distributed (iid), we expect a model trained on the target data ($D_{T}^{\prime}$) to perform the best on the test data $D_{T}^{\prime\prime}$. This performance value is the ceiling benchmark that we aim to match or surpass. For that purpose, we performed full fine-tuning (FFT) on all three base models using training data of all datasets (sentiment analysis and textual similarity), to come up with 18 FFT models for sentiment analysis and 9 FFT models for textual similarity. Hence, for any target dataset, we can have 15 DAFT models for sentiment analysis by excluding the three models that were fully fine-tuned on $D_{T}^{\prime}$, and 6 DAFT models for textual similarity task.
@@
We plot the zero-shot performance of model fine-tuned on one dataset and tested on another for both sentiment analysis and textual similarity tasks in Fig. 2 (b). It is important to note from the figure that most of the zero-shot performances are significantly better ($\gg 50\%$) than the performance of a model with untrained header layer ($50\%$ for binary classification). We also note that the performance of the Cornell and Tweet datasets do not match the other four datasets’ performance on the sentiment analysis task. Our empirical study suggests that the test data of Cornell is distributionally different from the other datasets, and for all FFT models the performance was poor. The Tweet dataset seems to have a different distribution in both its training and testing data (compared to the other 5 datasets). From the Tweet row and column, we observe that apart from the diagonal element, all the other accuracies are small, and a high $94\%$ accuracy is attained when the model is trained and tested on Tweet data. Upon inspecting the row of SST2, another interesting observation can be made, the performance of DAFT-Cornell beats that of DAFT-SST2. This is caused by the test dataset of SST2 being more aligned with the training data of Cornell (further details in Appendix A.3). For textual similarity datasets, the performance of DAFT models was not as good as the sentiment analysis, but still show promise. Lastly, to validate the transferability of the DAFT models to a new target domain or task, we looked at the LEEP scores (Nguyen et al., 2020) for all the 6 datasets of the sentiment analysis task (Fig. 2 (a)). Interestingly, the heat map with the LEEP scores closely follows the heat map that we generated using the performance of the DAFT models (Fig. 2 (b)). Thus, the transferability of DAFT models (corresponding LEEP scores) to a target domain is proportional to their (inference) performance on that target domain. These LEEP scores also tells us that a LEEP scored based ensemble method is possible and needs further investigation.
To demonstrate practical feasibility of DAFT models, we explored readily available DAFT models from public repositories. For target task we chose it to be question/answer in the field of Medical Knowledge and selected 9 sub-datasets from the MMLU dataset to be domain-adjacent of the target task (judging from the dataset names). These 9 sub-datasets are: Clinical Knowledge (CLNC KNW), College Biology (CLG BIO), Professional Medicine (PRF MED), Virology (VIROL), Nutrition (NUTR), Anatomy (ANATOM), College Medicine (CLG MED), High School Biology (HGH SCH BIO), and Medical Genetics (MED GEN). For each of these sub-datasets, we searched for corresponding fine-tuned models in public repositories. We found that LoRA-tuned DAFT models were available for all of them, with Llama 3.1 serving as the base model. After evaluating the performance of these DAFT models on all 9 sub-datasets, we created a heatmap following a similar method used for Fig. 2(b), and the result is shown in Fig. 4. In the heatmap, columns represent the publicly available (LoRA-tuned) DAFT models, and rows correspond to the individual sub-datasets. It is noticeable that these DAFT models exhibit highly variable performance across different sub-datasets; however, for any specific sub-dataset, the performance across different DAFT models does not vary significantly.
@@
[linecolor=black!0,backgroundcolor=orange!10]DAFT models can potentially provide competitive zero-shot performance on new tasks.
Figure 4: Performance of DAFTs on MMLU medical tasks: The 9 datasets are related to medical tasks, and the corresponding 9 DAFT models were obtained from Hugging Face.
3.2.1 Performance of DAFT ${}^{\text{\!\scriptsize Z}}$ (Zero Shot)
Fig. 3 shows the relative improvement $(RI)$ of all the DAFT models for sentiment analysis task when compared to the single best performing DAFT model. [^4] Thus, all the $RI$ values (corresponding to the other DAFTs) will be non-positive (zero for the best single DAFT). The relative improvement $(RI)$ of all the DAFT models for textual similarity task is provided in Appendix (Fig. 19). From Fig. 3 it is evident that there are some DAFT models that are not as good as the single best and for some specific choices (as for Tweet), the performance can be poor. On the other hand, for zero shot classification with Bart-large-mnli, Roberta-large-mnli, the performance were quite well. However, the zero-shot prompting using Opt-1.3b showed poor performance on most of the datasets. However, these (8 times) larger models (with zero shot classification or prompting) still fail to beat the single best DAFT in all cases, except Tweet. The low performance of DAFT models on Tweet is potentially because of the difference in the nature of this dataset (discussed previously in Section 3.2 and is supported by Fig. 2).
{mdframed}
[ align=left, linecolor=black!0,backgroundcolor=orange!10]Zero-shot performance of DAFT models can vary significantly, with some DAFT models even outperforming larger LLMs, making the choice of the DAFT model critical.
3.2.2 Few Shot Comparison of FT & DA(FT)2
Table 2: Average Relative Improvement of DA(FT)2 compared to FT (Top 6: sentiment analysis tasks. Bottom 3: textual similarity tasks): Positive values show DA(FT)2 improvement over FT on target (domain); larger values imply higher improvements. Please see Table 9 and 10 in Appendix for detailed results. We see that DA(FT)2 is significantly better than FT in few-shot problems. (The same RI values are plotted as line plot on the right to capture the trend with training data size.)
Target 2-shot 4-shot 8-shot 16-shot 32-shot 64-shot 128-shot